# Spatial Reasoning with Denoising Models

### v1

@@ 3.1 General Framework
Belief propagation.
The above paradigm is loosely inspired by belief propagation through Bayesian networks (Pearl, 1982), propagating information from observed to unobserved variables via sequential probabilistic inference. However, the set of assumptions is drastically reduced. It does not make any assumptions about distributions of $x_{i}^{0}$ and does not strictly require the Markov assumption about conditional independence between variables, as required for tractability in Bayesian networks. Instead, the network can (but not has to) be conditioned on all other variables, utilizing compression capabilities of neural networks.
@@ 3.2.1 Noise level Sampling
Diffusion Forcing (Chen et al., 2024a) suggests to independently sample a noise level vector $\mathbf{t}\sim\mathcal{U}([0,1]^{n})$, ensuring a uniform marginal distribution, $t_{i}\sim\mathcal{U}(0,1)$. However, this leads to the mean $\bar{t}$ following a Bates distribution, which is highly concentrated around $0.5$ – significantly different from the distribution encountered during inference (c.f. Fig. 4(a)). As a result, a trained model is undertrained for early (and late) inference steps, where most patches have noise levels close to $1$ (and 00), heavily reducing performance as demonstrated in Sec. 4.6.
@@
Figure 4: The left plot shows the joint distribution of noise levels $t_{i}$ and mean noise levels $\bar{t}$ for Diffusion Forcing training, while the right plot shows the same for our Uniform $\bar{t}$ approach. In Diffusion Forcing, the distribution of $t_{i}$ is correctly modeled but $\bar{t}$ is overly concentrated around $\bar{t}=0.5$, which does not reflect the distribution seen during inference. Our Uniform $\bar{t}$ addresses this issue, though it introduces some bias in the distribution of individual $t_{i}$ which we address by per-patch loss weighting.
@@ 3.3 Sampling
(b) Counting Polygons FFHQ dataset
Figure 5: Two further datasets. In addition to MNIST Sudoku, we introduce two further datasets. (a) The counting pixel dataset requires the model to generate images with an equal number of pixels per color. (b) The counting polygon dataset requires that all images contain two numbers: one for the number of polygons and one for the number of polygon edges.
@@ 4.1 Benchmark Datasets
We introduce three different datasets to quantify reasoning capabilities. They are aimed at different aspects to be tested. The MNIST Sudoku dataset captures complex (NP-hard) dependencies that need to be understood. The Counting Pixels dataset is an easier task that can be solved in a greedy fashion. Finally, we introduce the Counting Polygons FFHQ dataset, which moves closer to real-world images.
@@
The game Sudoku involves complex spatial dependencies, whereas the distribution of individual cells being MNIST numbers is simple to fit by a generative model. Note that the general task of solving Sudokus for a variable grid size is NP-complete and that depending on the level of masking, there can be multiple valid solutions resulting in uncertainty.
@@
Counting Polygons FFHQ Dataset
We also introduce a dataset in which each image contains a set of randomly positioned polygons and numbers (see Fig. 5(b)), where: The number of polygons, $N$, is randomly sampled from $\{1,\dots,9\}$, each polygon has $K$ vertices, sampled from $\{3,\dots,7\}$ and the numbers $N$ and $K$ are explicitly present in the image. The model’s task is to generate images following this rule and therefore to capture the relationship between the number of polygons, vertices, and the digits.
To introduce real-world complexity, we place these synthetic objects over backgrounds sampled from the FFHQ dataset. This forces the model to notice and learn meaningful spatial dependencies in a setting much closer to the real world. We train a ResNet classifier to estimate correctness. More details are given in Appendix E.
@@ 4.2 Experimental Setup
We compare SRMs in pixel space with standard denoising diffusion models that are equivalent in architecture and training up to the sampling of individual spatial variables, which are chosen to be image patches of task-specific sizes for our visual reasoning tasks. Since our MNIST Sudoku benchmark can be seen as an instance of inpainting, the baseline diffusion model is additionally conditioned on the incomplete Sudoku and its mask via concatenation in the input. Note that this is not necessary for SRMs, as already given variables are identified by the noise level zero. If not stated otherwise, we use stochastic sampling similar to DDPM with $1000$ steps (network evaluations) independent of the method of sequentialization during sampling. While SRMs are agnostic to different architectures, we use lightweight versions of widely established 2D UNets with spatial attention in low-resolution layers to avoid results being dominated by too extreme overparameterization. We describe the exact sampling process in Appendix A and implementation details in Appendix F.
@@ 4.3 MNIST Sudoku Results
Figure 7: Sequentialization with uncertainty-based order. We visualize the sampling process with the current sample $\mathbf{x}^{\mathbf{t}}$, noise level $\mathbf{t}$, estimated uncertainty (darker = lower) $\sigma_{\theta}(\mathbf{x}^{\mathbf{t}})$, and the single step to $\mathbf{t}=\mathbf{0}$ result $\hat{\mathbf{x}}^{\mathbf{0}}$. SRMs are able to reason over spatial variables by capturing complex dependencies (first two blocks) and handle uncertainty, e.g., in the case of multiple valid solutions as in the last block.
@@ 4.4 Even Pixels Results
4.5 Counting Polygon Results
For the final counting polygons on FFHQ dataset, we provide the quantitative results in Tab. 3. SRM outperforms the diffusion model consistently independent of the sampling method. This indicates potential of our training with individual noise levels, as this form of spatial disentanglement might be generally beneficial for certain data distributions. Unlike for the other two benchmarks, there is no clearly winning sampling strategy, as all methods perform similarly in terms of generating matching polygon and vertex numbers. We attribute this to two main differences of the dataset. First, using real images for backgrounds makes the task significantly more complex, as the denoising objective becomes less sensitive to the dependencies between numbers and polygons and capacity of the denoising network is spent for fitting the distribution of FFHQ faces. Secondly, the simultaneous generation of matching numbers and polygons can be approached in a coarse-to-fine manner with the numbers being more high-frequency details compared to larger low-frequency polygons. Still, the task of counting is inherently sequential and we can see the best result for a random order with overlapping patch denoising.
@@ 4.5 Counting Polygon Results
Diffusion Model 0.132
Ours, Parallel 0.166
Ours, Predicted Order w/o Overlap 0.144
Ours, Predicted Order + Overlap 0.142
Ours, Random Order w/o Overlap 0.164
Ours, Random Order + Overlap 0.186
Table 3: Counting Polygon FFHQ Experiments. Although this task is generally hard and not solved well, all SRM settings clearly outperform standard diffusion models.
@@ 4.6 Ablations
In Tab. 4, we provide the results of an ablation w.r.t. the noise level sampling from Sec. 3.2.1 on the hard difficulty of the MNIST Sudoku dataset. Enforcing a uniformly distributed mean during training is essential for all sampling methods. While independent uniform sampling of noise levels might be good enough for a very small number of variables as in Diffusion Forcing (Chen et al., 2024a), high and low mean noise levels quickly become severely undersampled when we increase it. As these two cases represent start and end points at test time, the input to the network is immediately out of the training distribution during sampling. We provide additional ablations in Appendix G.
@@ 5 Conclusion
We introduced Spatial Reasoning Models (SRMs), a framework that allows to reason over sets of continuous, spatial variables. Our framework allows parameterization for several different strategies of sequentialization and generation order. We introduced three benchmark tasks, which demonstrate that SRMs significantly outperform standard diffusion models in higher level reasoning tasks. Of particular interest is an automatic order prediction based on uncertainty.
@@
While we made significant steps forward, our models are still far away from solving the given tasks in an optimal fashion. However, we believe that the presented paradigm is able to achieve even better results. A particular topic for future research can be more advanced strategies for automatic prediction of generation order. Also, we speculate that backtracking-like strategies that allow to increase noise levels again hold much potential.
@@ Appendix D Even Pixels Evaluation
Appendix E Counting Polygons Dataset
@@
The color of the numbers and polygons is chosen in HSV space $\text{hsv}(c,1.0,0.9)$, where
@@ E.2 Evaluation
To assess the generative model’s ability to learn the implicit counting constraint, we employ a ResNet-50-based classifier with three independent prediction heads:
@@
  • Predicting the number of polygons.
  • Predicting the number of vertices per polygon.
The classifier is trained on a dataset with similar image compositions but without enforcing any consistency between the numbers and the actual object counts. For evaluation, the classifier detects the numbers present in the generated image and compares them to the true polygon and vertex counts. A generated image is classified as correct if the set of extracted numbers is the same as the set of actual polygon and vertex counts (regardless of order): $\text{correct}\Leftrightarrow\{\text{num}_{1},\text{num}_{2}\}=\{\#\text{polygons},\#\text{vertices}\}$. The evaluation metric accuracy is computed as the fraction of images on which this number-object correspondence is correct.
@@ Appendix F Implementation Details
For reasoning in the spatial visual domain, we divide images into patches of a fixed size. While we use the size of a cell, i.e., MNIST example $28\times 28$ as the patch size for the MNIST Sudoku dataset with image resolution $252\times 252$, we choose patch sizes of $4\times 4$ and $8\times 8$ for the even pixel and counting polygons datasets with resolutions $32\times 32$ and $128\times 128$, respectively.
@@
MNIST Sudoku Even Pixels Counting Polygons FFHQ
@@
Appendix G Additional Ablations
In Tab. 7, we ablate the choice of the sharpness hyperparameter in our noise level sampling algorithm for training SRMs. On the Even Pixels dataset, we can oversample noise level combinations for spatial variables that are more likely in sequential sampling strategies by lowering the sharpness. This can result in additional performance boosts depending on the data distribution.

### v2

@@ 3.1 General Framework
Belief Propagation
The above paradigm is loosely inspired by belief propagation through Probabilistic Graphical Models (Pearl, 1982), propagating information from observed to unobserved variables via sequential probabilistic inference. However, the set of assumptions is drastically reduced. It does not make any assumptions about distributions of $x_{i}^{0}$ and does not strictly require the Markov assumption about conditional independence between variables. Instead, the network can (but not has to) be conditioned on all other variables, utilizing compression capabilities of neural networks. Due to the nature of generative models, however, the formulation does not allow for explicit representation of probabilities but allows only sampling.
@@ 3.2.1 Noise Level Sampling
Diffusion Forcing (Chen et al., 2024a) suggests to independently sample a noise level vector $\mathbf{t}\sim\mathcal{U}([0,1]^{n})$, ensuring a uniform marginal distribution, $t_{i}\sim\mathcal{U}(0,1)$. However, this leads to the mean $\bar{t}$ following a Bates distribution, which is highly concentrated around $0.5$ – significantly different from the distribution encountered during inference (c.f. Fig. 4(a)). As a result, a trained model is undertrained for early (and late) inference steps, where most patches have noise levels close to $1$ (and 00), reducing performance (c.f. Sec. 4.6).
@@
Figure 4: Noise Level Sampling. We visualize the joint distribution of noise levels $t_{i}$ and mean noise levels $\bar{t}$ for (a) Diffusion Forcing training and (b) our Uniform $\bar{t}$ approach. In Diffusion Forcing, the distribution of $t_{i}$ is correctly modeled but $\bar{t}$ is overly concentrated around $\bar{t}=0.5$, which does not reflect the distribution seen during inference. Our Uniform $\bar{t}$ addresses this issue, though it introduces some bias in the distribution of individual $t_{i}$ which we address by per-patch loss weighting.
@@ 3.3 Sampling
(b) Counting Polygons / Stars FFHQ dataset
Figure 5: Two Further Datasets. In addition to MNIST Sudoku, we introduce two further datasets. (a) The Even Pixels dataset requires the model to generate images with an equal number of pixels per color. (b) Counting Polygons / Stars FFHQ requires that all images contain two digits: the number of objects and the number of polygon edges or star points, respectively. We introduce real-world complexity by placing the synthetic objects over FFHQ samples as backgrounds.
@@ 4.1 Benchmark Datasets
We introduce three different datasets to quantify reasoning capabilities. They are aimed at different aspects to be tested. The MNIST Sudoku dataset captures complex (NP-hard) dependencies that need to be understood. The Even Pixels dataset is an easier task that can be solved in a greedy fashion. Finally, we introduce the Counting Polygons / Stars FFHQ dataset, which moves closer to real-world images.
@@
The game Sudoku involves complex spatial dependencies, whereas the distribution of individual cells being MNIST numbers is simple to fit by a generative model. Note that the general task of solving Sudokus for a variable grid size is NP-complete and that depending on the level of masking, there can be multiple valid solutions resulting in uncertainty. We further highlight the difficulty of even a discrete Sudoku version for state-of-the-art LLMs in Appendix H.
@@
Counting Polygons / Stars FFHQ Dataset
We also introduce two datasets in which each image contains a set of randomly positioned polygons or stars and numbers (see Fig. 5(b)), where: The number of synthetic objects, $N$, is randomly sampled from $\{1,\dots,9\}$, each polygon / star has $K\in\{3,\dots,7\}$ vertices or $K\in\{2,\dots,9\}$ points, respectively, and the numbers $N$ and $K$ are explicitly present in the image. The model’s task is to generate images following this rule and hence to capture the relationship between the number of polygons / stars, vertices / points, and the digits.
To introduce real-world complexity, we place these synthetic objects over backgrounds sampled from the FFHQ dataset. This forces the model to notice and learn meaningful spatial dependencies in a setting much closer to the real world. We train ResNet classifiers to estimate correctness of generated samples and verify their capability on held-out validation splits. More details are given in Appendix E.
@@ 4.2 Experimental Setup
We compare SRMs in pixel space with standard denoising diffusion models that are equivalent in architecture and training up to the sampling of individual spatial variables, which are chosen to be image patches of task-specific sizes for our visual reasoning tasks. Since our MNIST Sudoku benchmark can be seen as an instance of inpainting, the baseline diffusion model is additionally conditioned on the incomplete Sudoku and its mask via concatenation in the input. Note that this is not necessary for SRMs, as already given variables are identified by the noise level zero. If not stated otherwise, we use stochastic sampling similar to DDPM with a total of $1000$ steps (network evaluations) independent of the method of sequentialization during sampling. Therefore, the computational cost is equal for all sampling methods resulting in a fair comparison. This also means that there are fewer denoising steps for individual variables the higher the degree of sequentialization (lower overlap). SRMs are agnostic to both noise schedules and denoiser architectures. For the main paper, we use rectified flows (Liu et al., 2023) and lightweight versions of widely established 2D UNets with spatial attention in low-resolution layers to avoid results being dominated by too extreme overparameterization. We report additional results with the cosine noise schedule (Nichol & Dhariwal, 2021) and Diffusion Transformers (DiT) (Peebles & Xie, 2023) in Appendix G, describe the exact sampling process in Appendix A, and provide implementation details in Appendix F.
@@ 4.3 MNIST Sudoku Results
Figure 7: Sequentialization With Predicted Order. We visualize the sampling process with the current sample $\mathbf{x}^{\mathbf{t}}$, noise level $\mathbf{t}$, estimated uncertainty (darker = lower) $\sigma_{\theta}(\mathbf{x}^{\mathbf{t}})$, and the single step to $\mathbf{t}=\mathbf{0}$ result $\hat{\mathbf{x}}^{\mathbf{0}}$. SRMs are able to reason over spatial variables by capturing complex dependencies (first two blocks) and handle uncertainty, e.g., in the case of multiple valid solutions as in the last block.
@@ 4.4 Even Pixels Results
4.5 Counting Polygons / Stars Results
For the final Counting Polygons / Stars on FFHQ datasets, we provide the quantitative results for sample accuracy (fulfillment of dataset rule) and consistency of objects (same number of polygon vertices / star points) in Tab. 3. SRM outperforms the diffusion model for almost all sampling methods. This indicates potential of our training with individual noise levels, as this form of spatial disentanglement might be generally beneficial for certain data distributions.
For the Counting Polygons dataset variant, there is no clearly winning sampling strategy, unlike for the other two benchmarks. We attribute this to two main differences of the dataset. First, using real images for backgrounds makes the task significantly more complex, as the denoising objective becomes less sensitive to the dependencies between numbers and polygons and capacity of the denoising network is spent for fitting the distribution of FFHQ faces. Secondly, the simultaneous generation of matching numbers and polygons can be approached in a coarse-to-fine manner with the numbers being more high-frequency details compared to larger low-frequency polygons. For MNIST Sudoku, all numbers have the same sizes such that, during sampling with a parallel denoising strategy, the commitment to individual digits has to happen at similar points in time. Due to the spatial dependencies of Sudoku, this is suboptimal and a spatially autoregressive strategy together with a good order can commit to the digits in a cell-by-cell fashion. For the Counting Polygons dataset, the numbers are high-frequency details compared to larger polygons. As a result, a coarse-to-fine generation with the diffusion baseline can first commit to a number of polygons of a certain vertex count and then generate matching digits.
We further validate this hypothesis by replacing the low-frequency polygons with stars that are composed of higher-frequencies, moving the “point of commitment” to numbers and stars closer together in time. For this dataset variant, we observe that the accuracy of the diffusion baseline decreases significantly compared to the version with polygons ($7\%$ vs $13.2\%$), while sequential sampling with overlap and predicted order maintains the same performance. More interestingly, for parallel generation, we noticed hallucinations of samples with stars having inconsistent numbers of points (cf. consistency column). For sequential sampling, the model can replicate stars after the first one has been generated, whereas in parallel sampling, the decisions over the number of points for all stars are again closer in time. We hypothesize that this behavior is not visible for polygons because of differences in terms of frequencies, with the diffusion model having more “time to correct itself” for low-frequency polygons, as their generation starts earlier than for the high-frequency stars.
@@ 4.5 Counting Polygons / Stars Results
Diffusion Model 0.132 0.990 0.070 0.544
Ours, Parallel 0.166 0.956 0.034 0.576
Ours, Predicted Order w/o Overlap 0.144 0.994 0.076 0.844
Ours, Predicted Order + Overlap 0.142 0.986 0.150 0.888
Ours, Random Order w/o Overlap 0.164 0.998 0.080 0.872
Ours, Random Order + Overlap 0.186 0.964 0.104 0.938
Table 3: Counting Polygons / Stars FFHQ Experiments. Although this task is generally hard and not solved well, sequentialization is beneficial, especially for consistency.
@@ 4.6 Ablations
In Tab. 4, we provide the results of an ablation of the noise level sampling from Sec. 3.2.1 on the hard difficulty of the MNIST Sudoku dataset. Enforcing a uniformly distributed mean during training is essential for all sampling methods. While independent uniform sampling of noise levels might be good enough for a very small number of variables as in Diffusion Forcing (Chen et al., 2024a), high and low mean noise levels quickly become severely undersampled when we increase it. As these two cases represent start and end points at test time, the input to the network is immediately out of the training distribution during sampling. We provide additional ablations w.r.t. stochasticity of inference, denoiser architecture, noise schedule, and the sharpness hyperparameter (cf. Sec. 3.2.1) in Appendix G.
@@ 5 Conclusion
We introduced Spatial Reasoning Models (SRMs), a framework that allows to reason over sets of continuous, spatial variables. Our framework allows parameterization for several different strategies of sequentialization and generation order. Of particular interest is an automatic order prediction based on uncertainty. The SRM framework also is agnostic to the employed denoising formulation and parameterization, and allows to use a wide variety of different denoising architectures. We introduced three benchmark tasks, which demonstrate that SRMs significantly outperform standard diffusion models in higher level reasoning tasks.
@@
While we made significant steps forward, our models are still far away from solving the given tasks in an optimal fashion. However, we believe that the presented paradigm is able to achieve even better results. A particular topic for future research can be more advanced strategies for automatic prediction of generation order. It is an interesting venture to further investigate what makes a distribution difficult to generate for a diffusion model. A potential difficulty metric could help optimizing for an optimal order. Also, we speculate that backtracking-like strategies that allow to increase noise levels again hold much potential.
@@ Appendix D Even Pixels Evaluation
Appendix E Counting Polygons/Stars Dataset
@@
The color of the numbers and polygons/stars is chosen in HSV space $\text{hsv}(c,1.0,0.9)$, where
@@ E.2 Evaluation
To assess the generative model’s ability to learn the implicit counting constraint, for each of the object types (polygons or stars) we employ a ResNet-50-based classifier with three prediction heads:
@@
  • Predicting the number of polygons/stars.
  • Predicting the number of object vertices.
For star objects, we only count the convex, outward-facing vertices. The classifiers are trained on a dataset with similar image compositions, but the consistency constraint between the numbers and the actual object counts is dropped in $50\%$ of cases. For all the prediction tasks mentioned above, both classifiers have achieved over $99.9\%$ accuracy in the validation set.
For evaluation, the classifier detects the numbers present in the generated image and compares them to the true object and vertex/spikes counts. A generated image is classified as correct if the set of extracted numbers is the same as the set of actual polygon/star objects and vertex counts (regardless of order): $\text{correct}\Leftrightarrow\{\text{num}_{1},\text{num}_{2}\}=\{\#\text{objects},\#\text{vertices}\}$.
The evaluation metric of the generative model, accuracy, is computed as the fraction of images for which the number-to-object-count correspondence is correct.
@@ Appendix F Implementation Details
For reasoning in the spatial visual domain, we divide images into patches of a fixed size. While we use the size of a cell, i.e., MNIST example $28\times 28$ as the patch size for the MNIST Sudoku dataset with image resolution $252\times 252$, we choose patch sizes of $4\times 4$ and $8\times 8$ for the even pixel and counting polygons/stars datasets with resolutions $32\times 32$ and $128\times 128$, respectively.
@@
MNIST Sudoku Even Pixels Counting Polygons/Stars FFHQ
@@
[…]
Easy Medium Hard
Metric Sampling UNet DiT UNet DiT UNet DiT
Accuracy $\uparrow$ Parallel 0.998 0.992 0.590 0.328 0.010 0.000
Predicted Order w/o Overlap 0.998 0.988 0.754 0.618 0.516 0.248
Predicted Order + Overlap 1.000 0.992 0.716 0.598 0.422 0.166
Random Order w/o Overlap 0.998 0.992 0.568 0.346 0.024 0.004
Random Order + Overlap 1.000 0.974 0.542 0.330 0.020 0.004
Graph-based Order w/o Overlap 0.996 0.978 0.662 0.470 0.266 0.086
Graph-based Order + Overlap 1.000 0.992 0.576 0.430 0.238 0.056
L1 Error $\downarrow$ Parallel 0.012 0.044 3.312 10.940 19.156 39.048
Predicted Order w/o Overlap 0.012 0.068 1.616 3.960 3.212 8.620
Predicted Order + Overlap 0.000 0.060 1.816 4.216 4.340 11.244
Random Order w/o Overlap 0.012 0.052 3.828 9.184 12.612 23.792
Random Order + Overlap 0.000 0.172 4.172 10.324 13.896 25.368
Graph-based Order w/o Overlap 0.020 0.180 2.544 5.864 5.816 12.752
Graph-based Order + Overlap 0.000 0.072 3.240 7.100 6.036 14.912
Table 8: Denoising Architecture Ablation on MNIST Sudoku. We compare SRM variantes using UNet and DiT denoising backbones. For each metric, model, and difficulty, the best-performing configuration is highlighted in bold. Sequential generation with predicted order yields the strongest performance across both metrics independent of the used architecture.
[…]
Easy Medium Hard
Metric Sampling Linear Cosine Linear Cosine Linear Cosine
Accuracy $\uparrow$ Parallel 0.998 1.000 0.590 0.622 0.010 0.018
Predicted Order w/o Overlap 0.998 1.000 0.754 0.748 0.516 0.378
Predicted Order + Overlap 1.000 0.998 0.716 0.706 0.422 0.298
Random Order w/o Overlap 0.998 1.000 0.568 0.502 0.024 0.036
Random Order + Overlap 1.000 0.994 0.542 0.510 0.020 0.024
Graph-based Order w/o Overlap 0.996 1.000 0.662 0.630 0.266 0.220
Graph-based Order + Overlap 1.000 0.996 0.576 0.620 0.238 0.210
L1 Error $\downarrow$ Parallel 0.012 0.000 3.312 3.048 19.156 17.140
Predicted Order w/o Overlap 0.012 0.000 1.616 1.812 3.212 5.140
Predicted Order + Overlap 0.000 0.012 1.816 2.056 4.340 5.920
Random Order w/o Overlap 0.012 0.000 3.828 4.812 12.612 13.232
Random Order + Overlap 0.000 0.032 4.172 5.020 13.896 14.200
Graph-based Order w/o Overlap 0.020 0.000 2.544 2.752 5.816 6.592
Graph-based Order + Overlap 0.000 0.032 3.240 2.924 6.036 6.848
Table 9: Noise Level Schedule Ablation on MNIST Sudoku. We compare SRM variants using Linear (Liu et al., 2023) and Cosine (Nichol & Dhariwal, 2021) noise level schedules. The winning schedule is highlighted in bold for each row. While the linear schedule clearly outperforms the cosine one for our setting, the benefits of sequentialization with a meaningful order remain independent of the choice of the noise level schedule.
Appendix G Additional Ablations
In Tab. 7, we ablate the choice of the sharpness hyperparameter in our noise level sampling algorithm for training SRMs. On the Even Pixels dataset, we can oversample noise level combinations for spatial variables that are more likely in sequential sampling strategies by lowering the sharpness. This results in higher performance depending on the data distribution.
To further demonstrate the architecture-agnostic nature of the SRM framework, we also trained a Diffusion Transformer (DiT B) (Peebles & Xie, 2023) with a patch size of 7 and 130M parameters on our MNIST Sudoku benchmark. The results, compared to those of a 2D UNet are presented in Tab. 8. Notably, in this highly structured domain, sequential generation with predicted variable order yields the best performance ndependent of the used architecture.
As an alternative to the linear schedule used in Rectified Flows (Liu et al., 2023), we also explored the cosine noise level schedule (Nichol & Dhariwal, 2021). The results are presented in Tab. 9. Independent of the noise level schedule, sequentialization with predicted order consistenly outperforms all other sampling strategies. Interestingly, the linear schedule seems to be more suitable than the cosine one for the dataset at hand. We suspect that this is because the cosine schedule allocates more time steps to low-noise regions, effectively focusing the denoising process on the high-frequency details, which are not relevant for the digit identity and thus not captured by our metrics.
Appendix H Sudoku Completion with LLMs
To evaluate the difficulty of generation-based Sudoku solving, we examined the performance of general-purpose large language models (LLMs). Specifically, we tested GPT-4o, GPT-4o-mini, and open-source Phi-4 on a text-based Sudoku completion task. Each model was given a textual representation of a Sudoku grid with missing elements and was tasked with filling in the blanks to complete the puzzle. The prompts included the full context of the Sudoku task: a description of the rules, output format, and three few-shot examples. An example is shown below:
[Example 1]
Input:
<sudoku>
34_|179|258
187|523|964
529|648|371
965|832|417
472|_16|835
_13|754|629
798|261|_43
631|485|792
_54|397|18_
</sudoku>
Output:
<sudoku>
346|179|258
187|523|964
529|648|371
965|832|417
472|916|835
813|754|629
798|261|543
631|485|792
254|397|186
</sudoku>
If a model’s output deviated from the required format, we resampled until a well-formed 9×9 grid of digits — suitable for evaluation — was produced. The model was then asked to provide only the full solution, with no ”inner thoughts”. This yielded an autoregressive generation variant similar to our image-based generation approach, but in the text domain.
Sudoku Difficulty
Method Modality Setup Easy Medium Hard
Diffusion Image Trained 0.994 0.536 0.008
SRM (Ours) 0.998 0.754 0.516
Phi-4 Text Few-shot 0.038 0.000 0.000
GPT-4o-mini 0.205 0.000 0.000
GPT-4o 0.556 0.001 0.011
Table 10: SRM vs LLM. To highlight the complexity of our MNIST Sudoku benchmark, we compare Sudoku solving capabilities of image generators with general-purpose LLMs used in a few-shot setting. Note that this is not a fair comparison, as SRM and Diffusion are trained directly for the task. Despite tackling a discrete and therefore simpler version of Sudoku, LLMs perform poorly in this task compared to SRM.
Table 10 compares the performance of the image-based Diffusion and SRM models with the few-shot LLMs across different difficulty levels. Although this comparison is not fair – our models are trained specifically for the task, in image space, whereas LLMs are few-shot prompted via text – it highlights the difficulty of solving Sudoku via generative models.
(a) Validity: proportion of outputs that are proper Sudoku grids (no duplicate digits).
(b) Consistency: solutions that match the initially masked grid entries.
(c) Accuracy: solutions that are both valid and consistent.
Figure 11: LLM Performance in Discrete Sudoku Completion. As the number of masked cells increases corresponding to an increasing difficulty level of the Sudoku completion task, both validity and consistency drop. Surprisingly, when almost all elements are masked, the accuracy improves – possibly due to similarity to some example grids in the prompt.
Figure 11 further breaks down model performance into validity, consistency, and overall accuracy as the number of masked cells increases. Validity measures whether the output follows the Sudoku rules, consistency ensures that the solution matches the initial grid, and accuracy requires both. Interestingly, while performance drops with increasing difficulty, accuracy improves when most of the board is masked, possibly due to resemblance with the few-shot examples in the prompt.
We additionally experimented with a chain-of-thought setup, where the model was prompted to iteratively fill one cell at a time, selecting its generation order. However, consistency with the input grid was insufficient for a reliable evaluation. We also tested the Deepseek-R1 reasoning model, allowing unconstrained internal reasoning before producing a final output. Although the model attempted to verify its own solutions, it frequently entered loops of self-correction or made mistakes while evaluating the constraints, resulting in unfinished rollouts or invalid Sudokus. Overall, the results suggest that solving Sudoku, even in simplified textual form, remains a challenging task for general-purpose generative models.